40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 20 12:55:02 2019, gautam, Update, CDS, c1iscaux testing 
    Reply  Mon Sep 23 10:49:34 2019, rana, Update, CDS, c1iscaux testing 
    Reply  Wed Sep 25 20:10:13 2019, Koji, Update, CDS, c1iscaux testing testDetectMons_190925.pngtestIPPOS_190925.png
       Reply  Thu Sep 26 20:09:40 2019, Koji, Update, CDS, c1iscaux testing Wht1.pdfWht2.pdfWht3.pdfWht4.pdflock.png
          Reply  Fri Sep 27 15:59:53 2019, gautam, Update, CDS, c1iscaux testing 
          Reply  Wed Oct 2 01:11:40 2019, Koji, Update, CDS, c1iscaux testing 8x
             Reply  Fri Oct 4 01:57:09 2019, Koji, Update, CDS, c1iscaux testing P_20191003_172956_vHDR_On.jpgTF.pdfPSD.pdf191003_AA_Filter.zip
                Reply  Sat Oct 5 00:03:21 2019, Koji, Update, CDS, c1iscaux testing REFL1_GAIN1.pdfREFL1_GAIN2.pdfREFL2_GAIN1.pdfREFL2_GAIN2.pdf
                   Reply  Tue Oct 8 03:32:42 2019, Koji, Update, CDS, CM servo board testing 7x
                      Reply  Tue Oct 8 17:59:29 2019, Koji, Update, CDS, CM servo board testing (portal) 6x
                         Reply  Tue Oct 8 18:42:39 2019, Koji, Update, CDS, CM servo board testing boosts.pdf
                            Reply  Mon Oct 14 16:06:28 2019, Koji, Update, CDS, CM servo board testing pole_zero_filter.pdf
                               Reply  Mon Oct 14 16:19:30 2019, Koji, Update, CDS, CM servo board testing testb2.pdf
                                  Reply  Mon Oct 14 16:25:03 2019, Koji, Update, CDS, CM servo board testing servo_out.pdf
                                     Reply  Mon Oct 14 16:34:42 2019, Koji, Update, CDS, CM servo board testing in12_output_offset.pdfin12_input_offset.pdfin12_input_offset2.pdf
                                        Reply  Mon Oct 14 17:32:28 2019, Koji, Update, CDS, Portal Elog entry for the recent CM servo board tests CM_Servo_Diagram.png
                         Reply  Tue Oct 8 20:23:03 2019, gautam, Update, CDS, c1iscaux testing 
    Reply  Mon Sep 30 11:20:43 2019, gautam, Update, CDS, c1iscaux testing - CM board code updated CMsoftTest.png
       Reply  Mon Sep 30 15:51:59 2019, gautam, Update, CDS, c1iscaux - some admin 
Message ID: 14906     Entry time: Wed Sep 25 20:10:13 2019     In reply to: 14903     Reply to this: 14908
Author: Koji 
Type: Update 
Category: CDS 
Subject: c1iscaux testing 

== Test Status ==
[none]
Whitening gain switching test
[none] AA enable/disable switching
[0th order] LO Det Mon channel check
[none] PD I/F board check
[done] QPD I/F board check
[none] CM Board
[none] ALS I/F board


- LO Det Mon channel check

The StripTool template for the test was made:
/cvs/cds/caltech/target/c1iscaux/testScripts/testDetectMons.str
Then, the RF output of the main Marconi was toggled a few times. -> Confirmed the channels are respopnding. (Attachment 1)

- IPPOS channel check

(0th order check) The StripTool template for the test was made:
/cvs/cds/caltech/target/c1iscaux/testScripts/testIPPOS.str
Then, the IPPOS QPD was shined with a phone LED. Initially I saw no response of the QPD. It turned out that the IPPOS IF module had no input cable connected. After the connection, all the 4 segments are responding to the phone LED and also the IFO beam.

(more careful check)
I decided to do more careful check of IPPOS. As there was a f~30mm lens on the oplev table, beam was focused such that only one element reacted to the incident beam. The beam power (a few mW) was too strong for a single QPD element, which saturates at ~6, an ND filter of OD0.6 was used to reduce the incident power.

Here are the results:
SEG1 (UPPER LEFT seen from the beam) | C1:ASC-IP_POS_QPD_Seg1_Mon 3.651+/-0.003 (N=10) | Incident Power 2.35+/-0.01 mW, QPD X_Calc (+) Y_Calc (+)

Segment Arrangement
(Seen from the beam)
Epics Channel CH output

Incident Power
(mW)

Polarity for the
X/Y_Calc channels
SEG1 UPPER LEFT C1:ASC-IP_POS_QPD_Seg1_Mon 3.651+/-0.003
(N=10)
2.35+/-0.01 X(+) / Y(+)
SEG2 LOWER LEFT C1:ASC-IP_POS_QPD_Seg2_Mon 3.607+/-0.002
(N=12)
2.35+/-0.01 X(+) / Y(-)
SEG3 LOWER RIGHT C1:ASC-IP_POS_QPD_Seg3_Mon 3.658+/-0.002
(N=11)
2.37+/-0.01 X(-) / Y(-)
SEG4 UPPER RIGHT C1:ASC-IP_POS_QPD_Seg4_Mon 3.529+/-0.004
(N=11)
2.30+/-0.01 X(-) / Y(+)

After the measurement, the lens and the filter were removed and the beam was adjusted to the center of the QPD.

Attachment 1: testDetectMons_190925.png  14 kB  Uploaded Wed Sep 25 21:14:33 2019  | Hide | Hide all
testDetectMons_190925.png
Attachment 2: testIPPOS_190925.png  20 kB  Uploaded Wed Sep 25 21:14:38 2019  | Hide | Hide all
testIPPOS_190925.png
ELOG V3.1.3-